OpenStudy (anonymous):

im helping my little brother with fraction & ive been out of school for a while & one of his question ask 2/6 of 18 pennies . Is the answer 1/6 ??

OpenStudy (phi):

you multiply 2/6 times 18/1 (top times top, bottom times bottom) 36/6 = 6

OpenStudy (phi):

Of course, if you notice 2/6 = 1/3 (divide top by 2, bottom by 2) then it is slightly simpler 1/3 * 18= 18/3 = 6 same answer (which is good)

OpenStudy (anonymous):

Ok Thank you Very Much !!!